Calculate the first cosmic velocity for the planet Jupiter. The radius of Jupiter is R = 7 × 10⁴ km, while the acceleration of free fall near the planet's surface is g = 26 m / s


1
Expert's answer
2021-01-29T20:02:08-0500

The first cosmic velocity is given as:


"v_1 = \sqrt{gR}"

where "g= 26m\/s^2" is the free fall nearJupiter's surface, and "R = 7\times 10^4km = 7\times 10^7m" is the radius of Jupiter. Thus, obtain:


"v_1 = \sqrt{26\cdot 7\times 10^7} \approx 42661.5m/s"

Answer. 42661.5 m/s.